OpenStudy (anonymous):

what is the x + 2 = 4 ???

OpenStudy (f_jayyy):

x=2

OpenStudy (anonymous):

what is 4- 2= ?

OpenStudy (f_jayyy):

x=4-2 x=2

OpenStudy (anonymous):

your supposed to try and let them figure it dude

OpenStudy (f_jayyy):

Sorry :) @TwentyOnePilots_

OpenStudy (chantysquirrel1129**):

You can report them c:

OpenStudy (chantysquirrel1129**):

Report them for giving direct answers :)

OpenStudy (charles.305):

I agree with Chanty. Your not supposed to give direct answers.